calculate volume of air needed for combustion of 1 kg of carbon
1
Expert's answer
2012-05-31T08:16:37-0400
The reaction is

C + O2 = CO2

The quantity of carbon is

n(C) = m(C)/M(C) = 1000g/12g/mol = 83.3 mol.

The quantity of O2 is

n(O2) = n(C) = 83.3 mol

The volume of O2 is

V(O2) = n(O2)*Vm = 83.3mol*22.4L/mol = 1866 L = 1.87 m^3

The volume of air is

V(air) = V(O2)/w(O2) = 1.87m^3/0.21 = 8.9 m^3